Show that the function f: [0,1]→ R defined by


F(x) = { 2 , when x is rational


{ 3 , when x is irrational


is not riemann integrable on[0,1]

1
Expert's answer
2022-03-01T13:36:46-0500

ANSWER.

Let "P=\left \{ 0=t_{0}<t_{1} <...<t_{n}=1\right \}" be a partition of the segment [0,1]. Denote "M_{k}=sup \left \{ f(t):t\in\left [t _{k-1},t_{k} \right ]\right \} ," "m_{k}=inf \left \{ f(t):t\in\left [t _{k-1},t_{k} \right ]\right \} , k=1,...,n" .

Since in any segment there are rational numbers and irrational numbers , then "M_{k}=3, m_{k}=2" for all "k=1,...,n" . So the upper Darboux sum "U(f,P)" respect to "P" is the sum

"U(f,P)=\sum _{1}^{n}M_{k} \left ( t_{k}-t_{k-1} \right )=\sum _{1}^{n}3\left ( t_{k}-t_{k-1} \right )=3(1-0)=3"

The lower Darboux sum "L(f,P)" is the sum

"L(f,P)=\sum _{1}^{n}m_{k} \left ( t_{k}-t_{k-1} \right )=\sum _{1}^{n}2\left ( t_{k}-t_{k-1} \right )=2(1-0)=2" .

Therefore, the upper Darboux integral

"U(f)= inf\left \{ U(f,P):P\, is \, \, partition \, \, of\, \, [0,1] \right \}=3,"

the lower Darboux integral

"L(f)= sup\left \{ L(f,P):P\, is \, \, partition \, \, of\, \, [0,1] \right \}=2."

The upper and lower Darboux integrals for "f" do not agree , so "f" is not integrable.